Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-price
Answer #1

Part A

ROI = (40% Margin × 0.6 Turnover) = 24.0%

0.6 Turnover = (Sales / $3000000 Average total assets)

Sales = 3000000*0.6 = $1800000

Part B

24.0% ROI = (30% Margin × Turnover)

Turnover = 0.80

0.80 Turnover = (Sales / $3000000 Average total assets)

Sales = $2400000

Part C

Increase in sales required = (2400000-1800000)/1800000 = 33.33%

Part D

Margin

Decreases

Turnover

Increases

ROI

Increases

Increase in marketing expenses, increases operating expenses and thus reduces margin.

Marketing efforts increases sales and thus there is increase in turnover. The increase in turnover is more than decrease in margin and thus there is increase in ROI.

Part E

Reduction in inventory carrying costs

Labor saving strategies

It is possible to reduce inventory carrying costs by limiting the inventory stock to those product lines generating the highest margins and/or highest turnover. This will reduce the amount of average total assets.

Sales force should be hired on commission basis, staff reduction during non-peak hours are some of the labor saving strategies that should be implemented.